# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 4.4.3 Getting one coordinate to dominate
$\mathbb{P}\left(\bigcap_{k=2}^{q}\left\{\bigcap_{i=0}^{L-1}\{\tau_{i+1,k}^{\prime}\leq\tau_{i,k}^{\prime}+n,\tilde{\tau}_{i+1,k}>\min\{\tau_{i,k}+n,\tau^{+}_{1},\tau_{L+1,k}\}\}\cap\{\tilde{\tau}_{L+1,k}>\tau_{L}+Ln\}\right\}\right)\geq 1-\sum_{i=0}^{L-1}3qr_{i}-o(1).$ (82)
@@
By Lemma 4.20, after $T=O(n\log n)$ steps, with probability $1-o_{\gamma}(1)$, the proportions chain has $S_{T,1}\geq S_{T,k}+\delta$ for all $k=2,...,q$. From there, Lemma 4.19 ensures that in a further $O(n)$ steps, the proportions vector is within an arbitrarily small $\rho$ distance from the fixed point corresponding to that phase, $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. Finally, Lemma 4.12 and Lemma 4.13 imply coupling of the proportions chain from there to that of a dynamics initialized from $\pi^{{\mathsf{ord}},1}$. ∎
@@
Suppose $\beta>\beta_{\mathsf{s}}$ and $\|S_{0}-(\frac{1}{q},...,\frac{1}{q})\|_{1}=O(n^{-1/2})$. Then for all $\gamma$ large, there exists a $C_{\gamma}>0$ such that the hitting time to having $S_{0,i}>S_{0,j}+\gamma n^{-1/2}$ for some $i$ and all $j\neq i$ is at most $C_{\gamma}n$ except with probability $1-o_{\gamma}(1)$.
@@
Applying Lemma 4.20, it is sufficient to show that in some $O(n)$ steps possibly depending on $\gamma$, the process attains $S_{0,1}>S_{0,i}+\gamma n^{-1/2}$ (up to permutation of the coordinates). In order to show this, we show that uniformly over any initialization, one obtains some coordinate which is $\gamma n^{-1/2}$ larger than the rest in some $n$ steps, with some $c_{\gamma}>0$ probability. This then gets boosted to $1-o_{\gamma}(1)$ probability after $C_{\gamma}n$ steps. There are two cases of initial proportions vector $S_{0}$ to consider:
@@
    The complement, in which case there are some $k\geq 2$ coordinates which are within $2\gamma n^{-1/2}$ of the maximal coordinate under the drift of (78) after $n$ steps.
In the first case, by Azuma–Hoeffding, after $n$ steps, the second term in (80) is at most $\gamma\sqrt{n}/C$ with probability $1-e^{-\Omega(\gamma^{2})}$, whence following the reasoning between (80)–(81), we have $\|S_{t,1}-\mathbf{S}_{t,1}\|\leq\gamma n^{-1/2}$ for all $t\leq n$ and we will have attained a configuration having $S_{T,1}>S_{T,i}+\gamma n^{-1/2}$ after $T=n$ steps.
@@
$S_{t,1}-\mathbf{S}_{t,1}=\sum_{l<t}(S_{t,1}-S_{t}-d_{\beta}(S_{t}))\,,$
to a random walk with a variance strictly bounded away from zero, taking $n$ steps that are $O(1/n)$ sized, there is a uniformly positive probability $p_{\gamma,1}$ that this is in $[2q\gamma n^{-1/2},3q\gamma n^{-1/2}]$. Also, conditionally on a typical realization of such a sequence, $S_{t,1}$ only changed $n/q+o(n)$ many times. On the remaining $(1-\frac{1}{q})n-o(n)$ steps, the increments of $S_{t,2}-\mathbf{S}_{t,2}$ still have a uniformly lower bounded variance. This leads to another uniformly positive probability $p_{\gamma,2}$ that it has $S_{t,2}-\mathbf{S}_{t,2}\in[\gamma n^{-1/2},-\gamma n^{-1/2}]$ (the conditioning on the event for $S_{t,1}$ only has a negligible effect on the drift for $S_{t,2}$ because of Lipschitz continuity of $d_{\beta}$). Repeating this for the next $q-1$ coordinates, with the very last one having no variance remaining but deterministically having decreased because of what happened for the other coordinates, we find that there is at least probability $p_{\gamma,1}\cdots p_{\gamma,q-1}$ such that
@@
4.5 Proof of Theorem 1.2
Proof of Theorem 1.2.
For item 1 of the theorem, fix $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, and suppose $m_{0}=m_{*}-\omega(n^{-1/2})$, where $m_{*}$ is specified in Lemma 4.1. If $\sigma_{0}$ is generated according to $\hat{\nu}^{\otimes}(m_{0})$, with probability $1-o(1)$, by concentration of multinomial random variables, $S_{0}$ has one coordinate that is within $O(n^{-1/2})$ of $m_{0}$ and the other coordinates are all within $O(n^{-1/2})$ of $\frac{1-m_{0}}{q-1}$. Without loss of generality, permute the coordinates so that it is the first coordinate that is close to $m_{0}$. Then, by Theorem 4.4, $\|\mathbb{P}_{\sigma_{0}}(S(\sigma_{T})\in\cdot)-\mathbb{P}_{\sigma_{0}^{\prime}}(S(\sigma_{T}^{\prime})\in\cdot)\|_{\textsc{tv}}=o(1)$, where $\sigma_{0}^{\prime}\sim\pi^{{\mathsf{dis}}}$. Furthermore, by Lemma 2.2, when $\beta<\beta_{\mathsf{c}}$, we have $\|\pi^{\mathsf{dis}}-\pi\|_{\textsc{tv}}=o(1)$, so it follows from the triangle inequality that
$\|\mathbb{P}_{\sigma_{0}\sim\hat{\nu}^{\otimes}(m_{0})}(S(\sigma_{T})\in\cdot)-\pi(S(\sigma)\in\cdot)\|_{\textsc{tv}}$ $\leq\|\mathbb{P}_{\sigma_{0}\sim\hat{\nu}^{\otimes}(m_{0})}(S(\sigma_{T})\in\cdot)-\mathbb{P}_{\sigma_{0}^{\prime}\sim\pi^{{\mathsf{dis}}}}(S(\sigma_{T}^{\prime})\in\cdot)\|_{\textsc{tv}}+\|\pi^{\mathsf{dis}}-\pi\|_{\textsc{tv}}$
$=o(1).$
Since $\hat{\nu}^{\otimes}(m_{0})$ and $\pi$ are invariant under permutation of vertices, we obtain that $\|\mathbb{P}_{\hat{\nu}^{\otimes}(m_{0})}(\sigma_{T}\in\cdot)-\pi\|_{TV}=o(1)$. For item 3, the proof also follows from Theorem 4.4 and Lemma 2.2 by a symmetrical argument, upon noticing that $\|\pi-\pi^{\mathsf{ord}}\|_{\textsc{tv}}=o(1)$, and $\pi^{{\mathsf{ord}}}$ is a $(1/q,...,1/q)$ mixture of $(\pi^{{\mathsf{ord}},i})_{i\in[q]}$, and each of the $q$ coordinates are equally likely under the initialization $\hat{\nu}^{\otimes}(m_{0})$ to dominate.
which will be less than $2\varepsilon$ for $\gamma$ large.
It remains to discuss Item 4. If $m_{0}\geq\frac{1}{q}+\omega(n^{-1/2})$, then together with spin symmetry and the vertex-permutation invariant of the initialization, Theorem 4.16 implies that $\|\mathbb{P}_{\hat{\nu}^{\otimes}(m_{0})}(\sigma_{T}\in\cdot)-\pi^{{\mathsf{ord}}}\|_{\textsc{tv}}=o(1)$ for $T=O(n\log n)$. If $\frac{1}{q}\leq m_{0}\leq\frac{1}{q}+\omega(n^{-1/2})$ then we first apply Lemma 4.21 to get a coordinate (by symmetry a uniform at random one) to obtain $\gamma n^{-1/2}$ separation from the rest, then apply Theorem 4.16 to get the same. On the other hand, by Lemma 2.2 we have $\|\pi^{\mathsf{ord}}-\pi\|_{\mathrm{TV}}=o(1)$, and thus item 4 follows from the triangle inequality. ∎
@@ 4.6 Lower bound on the mixing time with different choices of $m_{0}$
For every $q>2$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$, if $m_{*}(\beta,q)$ and $\hat{c}_{*}(q)$ are as in Theorem 1.2, then the CM dynamics initialized from $\hat{\nu}(m_{0})$ with
@@
We provide the details for the proof of item 1, the other cases following by similar reasoning. For any initialization parameter $m_{0}>m_{*}-Kn^{-1/2}$ for some $K=O(1)$, by Lemma 4.15 and 4.14, there is a positive probability $c_{K}>0$ that the process $S_{t,1}$ hits $m_{*}+\gamma n^{-1/2}$ before $m_{*}-\gamma n^{-1/2}$ (for sufficiently large $\gamma$) in some $t\leq C_{\gamma}n$ many steps. By Lemma 4.5, with probability $1-o(1)$, at exit, the configuration satisfies the necessary conditions to apply Theorem 4.4 and quasi-equilibrate to the ordered phase $\pi^{\mathsf{ord}}$. Putting these together, we find that for some $T_{0}=O(n\log n)$,
@@
Next, we claim that a Potts dynamics chain initialized from $\pi^{\mathsf{ord}}$ retains total-variation distance $1-o(1)$ for exponentially many steps to $\pi$ when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$. To see this, we use that by Corollary 2.2 the initialization from $\pi^{\mathsf{ord}}$ has proportions vector within $O(n^{-1/2})$ distance of $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$ (up to permutations). This is a stable fixed point of the drift function $d_{\beta}$ when $\beta>\beta_{\mathsf{u}}$ by Lemma 4.1. Following the Taylor expansion and martingale argument used in Lemma 4.5, in particular the application of Lemma 4.2, it takes exponential in $n$ time to leave an $\varepsilon$-neighborhood of the stable fixed point $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. (Note that the details of this last stage of reasoning are provided in the proof of [16, Theorem 3].) Combining the above, we find that for some $C>0$, for all $Cn\log n\leq t\leq e^{n/C}$,

### v2

@@ 4.4.3 Getting one coordinate to dominate
$\mathbb{P}$ $\left(\bigcap_{k=2}^{q}\left\{\bigcap_{i=0}^{L-1}\{\tau_{i+1,k}^{\prime}\leq\tau_{i,k}^{\prime}+n,\tilde{\tau}_{i+1,k}>\min\{\tau_{i,k}+n,\tau^{+}_{1},\tau_{L+1,k}\}\}\cap\{\tilde{\tau}_{L+1,k}>\tau_{L}+Ln\}\right\}\right)$
$\qquad\qquad\qquad\geq 1-\sum_{i=0}^{L-1}3qr_{i}-o(1)\,.$ (81)
@@
By Lemma 4.20, after $T=O(n\log n)$ steps, with probability $1-o_{\gamma}(1)$, the proportions chain has $S_{T,1}\geq S_{T,k}+\delta$ for all k=2,,qk=2,...,q. From there, Lemma 4.19 ensures that in a further $O(n)$ steps, the proportions vector is within an arbitrarily small $\rho$ distance from the fixed point corresponding to that phase, $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. Finally, Lemma 4.12 and Lemma 4.13 imply coupling of the proportions chain from there to that of a dynamics initialized from $\pi^{{\mathsf{ord}},1}$. ∎
@@
Suppose $\beta>\beta_{\mathsf{s}}$ and $\|S_{0}-(\frac{1}{q},...,\frac{1}{q})\|_{1}=O(n^{-1/2})$. Then for all $\gamma$ large, there exists a $C_{\gamma}>0$ such that the hitting time to having $S_{t,i}>S_{t,j}+\gamma n^{-1/2}$ for some $i$ and all $j\neq i$ is at most $C_{\gamma}n$ except with probability $1-o_{\gamma}(1)$.
@@
It is sufficient to show that in some $O(n)$ steps possibly depending on $\gamma$, the process attains $S_{t,1}>S_{t,j}+\gamma n^{-1/2}$ for all $j>1$ (up to permutation of the coordinates). In order to show this, we show that uniformly over any initialization, one obtains some coordinate which is $\gamma n^{-1/2}$ larger than the rest in $n$ steps, with some $c_{\gamma}>0$ probability. This then gets boosted to $1-o_{\gamma}(1)$ probability after $C_{\gamma}n$ steps. There are two cases of initial proportions vector $S_{0}$ to consider:
@@
    The complement, in which case there are at least 1 coordinate $i>1$ which is within $2\gamma n^{-1/2}$ of the maximal coordinate (assumed to be the first coordinate) under the drift of (77) after $n$ steps.
In the first case, by Azuma–Hoeffding’s inequality, after $n$ steps, the second term in the right-hand side of  (79) is at most $\gamma\sqrt{n}$ with probability $1-e^{-\Omega(\gamma^{2})}$, whence following the reasoning between (79)–(80), we have $\|S_{t}-\mathbf{S}_{t}\|_{1}\leq\gamma n^{-1/2}$ for all $t\leq n$ and we will have attained a configuration having $S_{T,1}>S_{T,i}+\gamma n^{-1/2}$ after $T=n$ steps.
@@
$S_{t,1}-\mathbf{S}_{t,1}=\sum_{l<t}(S_{l+1,1}-S_{l,1}-\frac{1}{n}d_{\beta}(S_{l}))\,$,
to a random walk with a variance strictly bounded away from zero, taking $n$ steps that are $O(1/n)$ sized, there is a uniformly positive probability $p_{\gamma,1}$ that this is in $2q\gamma n^{-1/2},3q\gamma n^{-1/2}]$. Also, conditionally on a typical realization of such a sequence, $S_{t,1}$ only changed $n/q+o(n)$ many times. On the remaining $(1-\frac{1}{q})n-o(n)$ steps, the increments of $S_{t,2}-\mathbf{S}_{t,2}$ still have a uniformly lower bounded variance. This leads to another uniformly positive probability $p_{\gamma,2}$ that it has $S_{t,2}-\mathbf{S}_{t,2}\in[-\gamma n^{-1/2},\gamma n^{-1/2}]$ (the conditioning on the event for $S_{t,1}$ only has a negligible effect on the drift for $S_{t,2}$ because of Lipschitz continuity of $d_{\beta}$). Repeating this for the next q1q-1 coordinates, with the very last one having no variance remaining but deterministically having decreased because of what happened for the other coordinates, we find that there is at least probability $p_{\gamma,1}\cdots p_{\gamma,q-1}$ such that
@@ 4.5 Lower bound on the mixing time with different choices of $m_{0}$
For every $q>2$ and $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$, if $m_{*}(\beta,q)$ and $\hat{c}_{*}(q)$ are as in Theorem 1.2, then the Potts Glauber dynamics initialized from $\hat{\nu}^{\otimes}(m_{0})$ with
@@
We provide the details for the proof of item 1, the other cases following by similar reasoning. For any initialization parameter $m_{0}>m_{*}-Kn^{-1/2}$ for some $K=O(1)$, by Lemma 4.15 and 4.14, there is a positive probability $c_{K}>0$ that the process $\{\max_{i}S_{t,i}\}$ hits $m_{*}+\gamma n^{-1/2}$ before $m_{*}-\gamma n^{-1/2}$ (for sufficiently large $\gamma$) in some $t\leq C_{\gamma}n$ many steps. By Lemma 4.5, with probability 1o(1)1-o(1), at exit, the configuration satisfies the necessary conditions to apply the second part of Theorem 4.4 and quasi-equilibrate to the ordered phase $\pi^{\mathsf{ord}}$. Putting these together, we find that for some $T_{0}=O(n\log n)$,
@@
Next, we claim that a Potts dynamics chain initialized from $\pi^{\mathsf{ord}}$ retains total-variation distance 1o(1)1-o(1) for exponentially many steps to $\pi$ when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$. To see this, we use that by Corollary 2.2 the initialization from $\pi^{\mathsf{ord}}$ has proportions vector within $O(n^{-1/2})$ distance of $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$ (up to permutations). This is a stable fixed point of the drift function $d_{\beta}$ when $\beta>\beta_{\mathsf{u}}$ by Lemma 4.1. Following the Taylor expansion and martingale argument used in Lemma 4.5, in particular the application of Lemma 4.2, it takes $\exp(\Omega(n))$ time to leave an $\varepsilon$-neighborhood of the stable fixed point $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. (Note that the details of this last stage of reasoning are provided in the proof of [17, Theorem 3].) Combining the above, we find that for some $C>0$, for all $T_{0}\leq t\leq e^{n/C}$,
@@
4.6 Proof of Theorem 1.2
Proof of Theorem 1.2.
For item 1 of the theorem, fix $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, and suppose $m_{0}=m_{*}-\omega(n^{-1/2})$, where $m_{*}$ is specified in Lemma 4.1. If $\sigma_{0}$ is generated according to $\hat{\nu}^{\otimes}(m_{0})$, with probability 1o(1)1-o(1), by concentration of multinomial random variables, $S_{0}$ has one coordinate that is within $O(n^{-1/2})$ of $m_{0}$ and the other coordinates are all within $O(n^{-1/2})$ of $\frac{1-m_{0}}{q-1}$. Without loss of generality, permute the coordinates so that it is the first coordinate that is close to $m_{0}$. Then, by Theorem 4.4, $\|\mathbb{P}_{\sigma_{0}}(S(\sigma_{T})\in\cdot)-\mathbb{P}_{\sigma_{0}^{\prime}}(S(\sigma_{T}^{\prime})\in\cdot)\|_{\textsc{tv}}=o(1)$, where $\sigma_{0}^{\prime}\sim\pi^{{\mathsf{dis}}}$. Furthermore, by Lemma 2.2, when $\beta<\beta_{\mathsf{c}}$, we have $\|\pi^{\mathsf{dis}}-\pi\|_{\textsc{tv}}=o(1)$, so it follows from the triangle inequality that
$\|\mathbb{P}_{\sigma_{0}\sim\hat{\nu}^{\otimes}(m_{0})}(S(\sigma_{T})\in\cdot)-\pi(S(\sigma)\in\cdot)\|_{\textsc{tv}}$
$\leq\|\mathbb{P}_{\sigma_{0}\sim\hat{\nu}^{\otimes}(m_{0})}(S(\sigma_{T})\in\cdot)-\mathbb{P}_{\sigma_{0}^{\prime}\sim\pi^{{\mathsf{dis}}}}(S(\sigma_{T}^{\prime})\in\cdot)\|_{\textsc{tv}}+\|\pi^{\mathsf{dis}}-\pi\|_{\textsc{tv}}$
$=o(1).$
Since $\hat{\nu}^{\otimes}(m_{0})$ and $\pi$ are invariant under permutation of vertices, we obtain that
$\|\mathbb{P}_{\hat{\nu}^{\otimes}(m_{0})}(\sigma_{T}\in\cdot)-\pi\|_{TV}=o(1).$
For item 3, the proof also follows from Theorem 4.4 and Lemma 2.2 by a symmetrical argument, upon noticing that $\|\pi-\pi^{\mathsf{ord}}\|_{\textsc{tv}}=o(1)$, and $\pi^{{\mathsf{ord}}}$ is a (1/q,,1/q)(1/q,...,1/q) mixture of $(\pi^{{\mathsf{ord}},i})_{i\in[q]}$, and each of the $q$ coordinates are equally likely under the initialization $\hat{\nu}^{\otimes}(m_{0})$ to dominate.
which will be less than $2\varepsilon$ for $\gamma$ large.
It remains to discuss Item 4. If $m_{0}\geq\frac{1}{q}+\omega(n^{-1/2})$, then together with spin symmetry and the vertex-permutation invariant of the initialization, Theorem 4.16 implies that $\|\mathbb{P}_{\hat{\nu}^{\otimes}(m_{0})}(\sigma_{T}\in\cdot)-\pi^{{\mathsf{ord}}}\|_{\textsc{tv}}=o(1)$ for $T=O(n\log n)$. If $\frac{1}{q}\leq m_{0}\leq\frac{1}{q}+\omega(n^{-1/2})$ then we first apply Lemma 4.21 to get a coordinate (by symmetry a uniform at random one) to obtain $\gamma n^{-1/2}$ separation from the rest, then apply Theorem 4.16 to get the same. On the other hand, by Lemma 2.2 we have $\|\pi^{\mathsf{ord}}-\pi\|_{\mathrm{TV}}=o(1)$, and thus item 4 follows from the triangle inequality.
The slow mixing results are exactly the statement of Theorem 4.22. ∎
We also include a proof of Theorem 1.4 on initializations from the Potts Gibbs measure at a different temperature, to justify why it follows from the above arguments.
Proof of Theorem 1.4.
It is evident at this point that the only properties of the initialization used in our proof of Theorem 1.2 were on the size of its largest color class relative to $m_{*}n$, and the differences between its non-dominant color counts being $O(\sqrt{n})$. The requisite properties for samples from $\pi_{\beta_{0}}$ are given by Corollary 2.2. For the choice of $d_{*}(\beta,q)$, in item 1 when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{c}})$, we choose $d_{*}(\beta,q)=\beta_{\mathsf{c}}$, and in item 3, the choice is $d_{*}(\beta,q)=\inf\{d:\frac{\theta_{\mathsf{r}}(d)(q-1)+1}{q}>m_{*}(\beta)\}$, as the typical size of the largest coordinate in a sample from $\pi_{d}$ is $\frac{1}{q}(\theta_{\mathsf{r}}(d)(q-1)+1)$ per Corollary 2.2. Finally note that any $\pi_{\beta_{0}}$ could work for item 4 since the largest coordinate of a sample is always at least $1/q$. ∎